### v1

@@
Abstract: This paper investigates the structure of $K_{2}$-groups for certain quotient rings of the integral group ring $\mathbb{Z}[G]$. Let $G$ be a finite abelian $p$-group with $p$-rank $r>1$, let $\Gamma$ be the maximal $\mathbb{Z}$-order of $\mathbb{Q}[G]$, and let $\widetilde{G}$ denote the sum of all elements of $G$ in the group ring. By employing the framework of Kähler differentials, we first determine that the relative $K$-group $)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)) is an elementary abelian $p$-group of rank $r$. Building upon this result, we establish an explicit isomorphism:
@@
Key words: $K_{2}$-groups; group rings; finite abelian $p$-groups; Kähler differentials
@@
This paper is a continuation of our previous work [1]. Let $G$ be a finite abelian $p$-group with $p$-rank $r>1$, and let $\Gamma$ denote the maximal $\mathbb{Z}$-order in the group algebra $\mathbb{Q}[G]$. Set $J=|G|\Gamma$. By Proposition 2.2 in [2], $\Gamma$ is isomorphic to a direct product of rings of algebraic integers, with the inclusions $J\subseteq\mathbb{Z}[G]\subseteq\Gamma$ holding.
@@
In the present paper, we establish an explicit relation between the $K_{2}$-groups of the quotient rings appearing in this square. The paper is organized as follows: Section 2 collects the necessary preliminaries.
In Section 3, we utilize the framework of Kähler differentials to prove two key results, which are then applied to show that $)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)) is an elementary abelian $p$-group of rank $r$ (see Lemma 3.1). We also demonstrate that the following exact sequence of finite abelian $p$-groups splits:
@@
Furthermore, Theorem 2.3 is employed to establish the excision property associated with (0.1), as detailed in Proposition 3.2. As a consequence, we obtain the isomorphism:
$).K_{2}(\mathbb{Z}[G]/I,J/I)\cong K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)).
Finally, by extending the methods of [3] to a more general setting, we prove our main result, Theorem 3.4, in the last section. We show that the exact sequence of finite abelian $p$-groups
also splits. Combining this with the aforementioned excision isomorphism, we establish the decomposition:
@@
This result provides a structural decomposition for the $K_{2}$-groups of the quotient rings in square (0.1), successfully extending our previous computational studies to a general theoretical framework.
@@
$I/I^{2}\xrightarrow{d}T\otimes_{S/R}\Omega_{S/R}\xrightarrow{\mathrm{D}\pi}\Omega_{T/R}\rightarrow 0,$
@@
For $s,t\in A$, let $\alpha_{0}=s$ and $\alpha_{i}=b_{i}$ for $1\leq i\leq r$. By applying Lemma 1.2 to the first component of the symbol $\langle b_{1}\cdots b_{r}s,b_{1}\cdots b_{r}t\rangle$, we have:
@@
$=1.$ (as $b_{i}J\subset I$)
Since $r>1$ and $b_{i}J\subset I$ for each $i$, the elements $(b_{1}\cdots b_{r})^{2}t$ and $(b_{1}\cdots b_{r})^{2}st/b_{i}$ are all contained in $I$. In the relative group $K_{2}(A/I,J/I)$, these elements represent $0\in A/I$, which implies each Dennis-Stein symbol in the product evaluates to the identity $1$. Thus, $\operatorname{Im}(\psi)=1$, it follows that $K_{2}(A/I,J/I)\cong J/I\otimes_{A/J}\Omega_{A/J}$.
To complete the proof, it suffices to show that $J/I\otimes_{A/J}\Omega_{A/J}\cong J/I\otimes_{A/I}\Omega_{A/I}$. According to Lemma 1.3, there is an exact sequence of $A/J$-modules:
@@
In this section, we extend several results from [3, Section 4].
@@
Let $G$ be as above with $|G|>2$, and set $x_{i}=g_{i}-1$. Then $)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)) is an elementary abelian $p$-group of rank $r$ with the basis $\{\langle x_{i},\prod_{j=1}^{r}x_{j}^{p^{n_{j}}-1}\rangle|1\leq i\leq r\}.$
@@
According to Theorem 2.3, there exists an isomorphism
@@
By Lemma 2.1, $\Omega_{\mathbb{F}_{p}[G]}$ is a free $\mathbb{F}_{p}[G]$-module of rank $r$. Since $p\widetilde{G}=0$ in $\mathbb{F}_{p}[G]$, the group $)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)) has exponent $p$, which implies that it is an elementary abelian $p$-group of rank $r$. Note that within $\mathbb{F}_{p}[G]$, we have the following identity:
Consequently, by Lemma 2.1, the set $\{(\prod_{j=1}^{r}x_{j}^{p^{n_{j}}-1})\otimes\mathrm{d}x_{i}\mid 1\leq i\leq r\}$ constitutes a basis for $(\widetilde{G})\otimes_{\mathbb{F}_{p}[G]}\Omega_{\mathbb{F}_{p}[G]}$. Via the map $\rho$ defined in Theorem 2.3, this basis corresponds to the generators of $)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)), and the result follows.
@@
Let $G,I,J$ as defined above with $|G|>2$. Then
@@
Note that since $I_{1}=J_{1}\cap I_{2}$, we have $(J_{1}/I_{1})\cap(I_{2}/I_{1})=0$. This allows for an alternative proof of the isomorphism. By [9, Theorem 14], there exists an exact sequence:
$(J_{1}/I_{1})/(J_{1}/I_{1})^{2}\otimes_{A^{e}}(I_{2}/I_{1})/(I_{2}/I_{1})^{2}\xrightarrow{\psi}K_{2}(A/I_{1},J_{1}/I_{1})\rightarrow K_{2}(A/I_{1},J_{2}/I_{2})\rightarrow 0,$
where $A^{e}=A\otimes_{\mathbb{Z}}A^{op}$ and the map $\psi$ is defined by $\psi(\overline{a}\otimes\overline{b})=\langle a,b\rangle$. In our case, since $(J_{1}/I_{1})^{2}=0$, Remark 1.1 ensures that the Dennis-Stein symbol $\langle a,b\rangle$ is additive in both components. Given that $p\widetilde{G}\in(J_{1}/I_{1})\cap(I_{2}/I_{1})=0$ in $A/I_{1}$, for any $s,t\in A$, we have:
$\psi(\overline{\widetilde{G}s}\otimes\overline{pt})=\langle\widetilde{G}s,pt\rangle=\langle p\widetilde{G}s,t\rangle=\langle 0,t\rangle=1.$
Hence, the image of $\psi$ is trivial, and the isomorphism is established.
@@
Let $G,I,J$ as defined above with $r>1$. Then
[…]
@@
where the surjectivity of the maps follows from [2, Proposition 1.1]. The homomorphism $i$ is injective: each generator of $)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)) has order $p$, and each corresponding image of $i$ denoted by the same literal, also has order $p$ and lies in a basis for $K_{2}\left(\mathbb{F}_{p}[G]\right)$ [10]. As a result, the bottom exact sequence splits: the monomorphism $i$ has a section, which we denote by $j$. Since $g_{1}$ is an isomorphism by Proposition 3.2, the commutativity of the first square implies that $f_{1}$ is injective. Consequently, $g_{1}^{-1}jg_{2}$ is a section of $f_{1}$, showing that $f_{1}$ is a split injection. The top exact sequence therefore splits, and we obtain
[…]
@@
By Snake Lemma, we have $\mbox{Ker}(g_{2})\cong\mbox{Ker}(g_{3})$. Besides, the corresponding square of (0.1) for $K_{2}$-groups is also Cartesian.

### v2

@@
Abstract: This paper investigates the structure of $K_{2}$-groups for certain quotient rings of the integral group ring $\mathbb{Z}[G]$. Let $G$ be a finite abelian $p$-group with $p$-rank $r$, let $\Gamma$ be the maximal $\mathbb{Z}$-order of $\mathbb{Q}[G]$, and let $\widetilde{G}$ denote the sum of all elements of $G$ in the group ring. By employing the framework of Kähler differentials, we first determine that the relative $K$-group $)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)) is an elementary abelian $p$-group of rank $r$ when $|G|>2$. Building upon this result, we establish an explicit isomorphism for $r>1$:
@@
Key words: $K_{2}$-groups; group rings; Dennis-Stein symbols; Kähler differentials
@@
This paper is a continuation of our previous work [1]. Let $G$ be a finite abelian $p$-group with $p$-rank $r$, and let $\Gamma$ denote the maximal $\mathbb{Z}$-order in the group algebra $\mathbb{Q}[G]$. Set $J=|G|\Gamma$. By Proposition 2.2 in [2], $\Gamma$ is isomorphic to a direct product of rings of algebraic integers, with the inclusions $J\subseteq\mathbb{Z}[G]\subseteq\Gamma$ holding.
@@
In the present paper, we establish an explicit relation between the $K_{2}$-groups of the rings lying in (0.1). The rest of the article is organized as follows: Section 1 is devoted to the necessary preliminaries.
In Section 2, we employ the framework of Kähler differentials to prove two key results for $|G|>2$, which are then applied to determine the explicit structure of $)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)) (see Lemma 3.1). Furthermore, Theorem 2.3 is used to establish the excision property for $K_{2}$ with respect to the cartesian square (0.1), as described in Proposition 3.3.
As a consequence, in Section 3, we show that the following exact sequences of finite abelian $p$-groups split when $r>1$:
@@
Combining this with the aforementioned excision isomorphism, we establish the decomposition (see Theorem 3.5):
@@
This result, obtained by extending the methods of [3] to a more general setting, provides a structural decomposition for the $K_{2}$-groups of the quotient rings of $\mathbb{Z}[G]$ in square (0.1), and extends our previous computational studies to a general theoretical framework.
@@
$I/I^{2}\xrightarrow{\mathrm{d}}T\otimes_{S/R}\Omega_{S/R}\xrightarrow{\mathrm{D}\pi}\Omega_{T/R}\rightarrow 0,$
@@
For $s,t\in A$, let $\alpha_{0}=s$ and $\alpha_{i}=b_{i}$ for $1\leq i\leq r$. We have
@@
$=1.$ (since $b_{i}J\subset I$)
The second-last equality follows by regarding the first entry of $\langle b_{1}\cdots b_{r}s,b_{1}\cdots b_{r}t\rangle$ as a product of $r+1$ factors and leaving the second entry unchanged. The last equality holds because $r>1$ and $b_{i}J\subset I$ for each $i$, so that the elements $(b_{1}\cdots b_{r})^{2}t$ and $(b_{1}\cdots b_{r})^{2}st/b_{i}$ all lie in $I$. In the relative group $K_{2}(A/I,J/I)$, these elements represent 0 in $A/I$, which implies that every Dennis-Stein symbol in the product evaluates to the identity element $1$.
Thus, $\operatorname{Im}(\psi)=1$, it follows that $K_{2}(A/I,J/I)\cong J/I\otimes_{A/J}\Omega_{A/J}$. To complete the proof, it suffices to show that $J/I\otimes_{A/J}\Omega_{A/J}\cong J/I\otimes_{A/I}\Omega_{A/I}$. According to Lemma 1.3, there is an exact sequence of $A/J$-modules:
@@
In this section, we extend several results from [3, Section 4] concerning $K_{2}$-relations for rings lying in the Cartesian square (0.1).
@@
Let $G$ be as above with $|G|>2$, and set $x_{i}=g_{i}-1$. Then $)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)) is an elementary abelian $p$-group of rank $r$ with basis $\{\langle x_{i},\prod_{j=1}^{r}x_{j}^{p^{n_{j}}-1}\rangle\mid 1\leq i\leq r\}.$
@@
Note that $x_{i}G=0$ for all $1\leq i\leq r$, and the following identity holds in $\mathbb{F}_{p}[G]$:
Thus, $\widetilde{G}$ has degree greater than 1 if and only if $|G|>2$. Since $|G|>2$ holds, according to Theorem 2.3, there exists an isomorphism
@@
By Lemma 2.1, $\Omega_{\mathbb{F}_{p}[G]}$ is a free $\mathbb{F}_{p}[G]$-module of rank $r$. Since $p\widetilde{G}=0$ in $\mathbb{F}_{p}[G]$, the group $)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)) has exponent $p$, which implies that it is an elementary abelian $p$-group of rank $r$. Consequently, by Lemma 2.1, the set $\{(\prod_{j=1}^{r}x_{j}^{p^{n_{j}}-1})\otimes\mathrm{d}x_{i}\mid 1\leq i\leq r\}$ constitutes a basis for $(\widetilde{G})\otimes_{\mathbb{F}_{p}[G]}\Omega_{\mathbb{F}_{p}[G]}$. Via the map $\rho$ defined in Theorem 2.3, this basis corresponds to a basis of $)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)), and the result follows.
Remark 3.2.
$)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)) is trivial when $G=C_{2}$, as noted in [3, Lemma 4.1].
@@
Let $G,I,J$ be as above with $|G|>2$. Then
@@
Since $I_{1}=J_{1}\cap I_{2}$, we have $(J_{1}/I_{1})\cap(I_{2}/I_{1})=0$. Using this fact, by Theorem 14 in [9], we obtain an exact sequence:
$(J_{1}/I_{1})/(J_{1}/I_{1})^{2}\otimes_{A^{e}}(I_{2}/I_{1})/(I_{2}/I_{1})^{2}\xrightarrow{\psi}K_{2}(A/I_{1},J_{1}/I_{1})\rightarrow K_{2}(A/I_{1},J_{2}/I_{2})\rightarrow 1,$
where $A^{e}=A\otimes_{\mathbb{Z}}A^{op}$ and the map $\psi$ is defined by $\psi(\overline{a}\otimes_{A^{e}}\overline{b})=\langle a,b\rangle$. In our case, since $(J_{1}/I_{1})^{2}=0$, Remark 1.1 ensures that the Dennis-Stein symbol $\langle a,b\rangle$ is additive in both components. Given that $p\widetilde{G}\in(J_{1}/I_{1})\cap(I_{2}/I_{1})=0$ in $A/I_{1}$, for any $s,t\in A$, we have:
$\psi(\overline{\widetilde{G}s}\otimes_{A^{e}}\overline{pt})=\langle\widetilde{G}s,pt\rangle=\langle p\widetilde{G}s,t\rangle=\langle 0,t\rangle=1.$
Hence, the image of $\psi$ is trivial. This yields an alternative proof of the isomorphism.
@@
Let $G,I,J$ be as above with $r>1$. Then
$).K_{2}(\mathbb{Z}[G]/I)\cong K_{2}(\mathbb{Z}[G]/J)\oplus K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)).
@@
where the surjectivity of the maps follows from [2, Proposition 1.1]. The homomorphism $i$ is injective: each generator of $)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)) has order $p$, and each corresponding image of $i$ denoted by the same literal, also has order $p$ and lies in a basis for $K_{2}\left(\mathbb{F}_{p}[G]\right)$ when $r>1$ (see [10]). As a result, the bottom exact sequence splits: the monomorphism $i$ has a section, which we denote by $j$. Since $g_{1}$ is an isomorphism by Proposition 3.3, the commutativity of the first square implies that $f_{1}$ is injective. Consequently, $g_{1}^{-1}jg_{2}$ is a section of $f_{1}$, showing that $f_{1}$ is a split injection. The top exact sequence therefore splits, and we obtain
$).K_{2}(\mathbb{Z}[G]/I)\cong K_{2}(\mathbb{Z}[G]/J)\oplus K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)).
@@
By the Snake Lemma, we have $\mathrm{Ker}(g_{2})\cong\mathrm{Ker}(g_{3})$. In addition, the corresponding square for $K_{2}$-groups in (0.1) is Cartesian in the category of groups.